If I initially have a gas at a pressure of 15 atm, volume of 20 L, and a temperature of 50 k, and then I raise the volume and temperature of the gas to 45 L and 130 k, what is the pressure


1
Expert's answer
2021-02-28T06:18:57-0500

P*V/T = Const. P1*V1/T1 = P2*V2/T2, P2 = P1*V1/T1*T2/V2;

P2 = 15 atm*20/50*130/45 = 17.33 atm.